Line with slope 1/5 passes through (6,3). If the ordinate of a point on the line is 4, what is its abscissa?
Answer by TimothyLamb(4379)

y - y1 = m(x - x1)
y - 3 = (1/5)(x - 6)
y - 3 = (1/5)x - (1/5)6
y = (1/5)x - 6/5 + 15/5
y = (1/5)x + 9/5
---
given (x,y)
x = abscissa
y = ordinate
---
4 = (1/5)x + 9/5
20/5 - 9/5 = (1/5)x
11/5 = (1/5)x
---
Answer:
x = 11
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php